The [α] of pure quinine, an antimalarial drug, is −165. If a solution contains 79% quinine and 21% of its enantiomer (ee = 58%), what is [α] for the solution?
Answer #1

net rotation = 0.79-0.21=0.58

therefore [a] for solution = -0.58*165 = -95.7

So required [α] for the solution is -95.7